A particle is moving at a speed of 12m/s on a circle of radius r.If the speed of the particle is increased to 15m/s in 2 sec,what is the angular acceleration produced?
 

Grade:11

1 Answers

Khimraj
5 years ago
w1 = v1/r = 12/r
w2 = v2/r = 15/r
angular acceleration = (w2 – w1)/t = 3/2r
If value of radius is given then put in place of r
Hope it clears...................